### v1

# Reducing Vision Transformer Latency on Edge Devices via GPU Tail Effect and Training-free Token Pruning

Abstract
This paper investigates how to efficiently deploy transformer-based neural networks on edge devices. Recent methods reduce the latency of transformer neural networks by removing or merging tokens, with small accuracy degradation. However, these methods are not designed with edge device deployment in mind, and do not leverage information about the hardware characteristics to improve efficiency. First, we show that the relationship between latency and workload size is governed by the GPU tail-effect. This relationship is used to create a token pruning schedule tailored for a pre-trained model and device pair. Second, we demonstrate a training-free token pruning method utilizing this relationship. This method achieves accuracy-latency trade-offs in a hardware aware manner. We show that for single batch inference, other methods may actually increase latency by 18.6-30.3% with respect to baseline, while we can reduce it by 9%. For similar latency (within 5.2%) across devices we achieve 78.6%-84.5% ImageNet1K accuracy, while the state-of-the-art, Token Merging, achieves 45.8%-85.4%.
1 Introduction
In the past decade, Internet of Things (IoT) and commodity edge devices for various applications have become ubiquitous [4, 19, 38]. Edge devices have become sufficiently powerful, and model miniaturization techniques sufficiently capable, that machine learning models can be deployed to the network edge for various tasks, including computer vision applications [11, 32]. However, state of the art performance on various computer vision tasks are often claimed by Vision Transformer (ViT) based neural network [10] architectures. ViT models such as DeiT [34] or backbones such as DINOv2 [25] are not designed nor miniaturized for edge deployment.
Prior work [20] has shown that ViT’s feature high levels of redundancy that can be exploited for latency reduction benefits. One general approach is to identify parts of the input (tokens) which contain low information content and remove them. This is referred to as token sparsification. Training-free token sparsification methods such as Token Merging (ToMe) [1] have been effective at reducing latency of pre-trained models. Other approaches like DynamicViT [30] can yield better accuracy than training-free methods, but at the cost of training time on server-grade hardware.
Figure 1: Forward pass latency for widely used DeiT-B and DinoV2 models across various hardware (See Tab. 3) evaluated on the ImageNet1K [8] dataset. These plots demonstrate the relationship between workload size and latency can be variable and non-linear across a variety of hardware. Consequently, it is possible to achieve large latency reductions without removing too many tokens in many cases. Note that the latency axis is not necessarily indexed from 0 for all plots.
We address three shortcomings of existing techniques. (1) Many existing efficient methods do not consider fine-grained hardware or latency characteristics [1, 16, 37]. Fig. 1 demonstrates the diversity of latency-workload trends across devices and workload sizes. As a result, there is room to optimize efficient methods in a hardware aware manner by considering this relationship. (2) Some existing efficient methods may require extensive training [42, 30, 5] hinders the deployment of pre-trained models on edge devices. (3) Prior work has investigated hardware aware methods for CNN’s [41, 40], the GPU tail effect has not been investigated specifically utilized for transformer models. Given these properties of existing token removal methods, this paper focuses on reducing ViT latency by considering hardware characteristics and without requiring training. This paper has the following contributions:
  1. 1.
    Introduce the first method for deciding a hardware-aware ViT token pruning schedule for model and device pairs.
  2. 2.
    Develop a novel training-free token pruning mechanism. For similar latency (within 5.2%) across various hardware and workload sizes, we achieve 0.46-43.7 higher accuracy than ToMe, a popular state-of-the-art method.
2 Background and Related Work
In this section, we review related work on model acceleration and efficient methods for Vision Transformer (ViT) models. Some post-processing or fine-tuning methods such as quantization [18] are compatible with token sparsification techniques such as our method.
2.1 Model Acceleration of Vision Transformers
ViT  [10] architectures such as DINOv2 [25] have achieved state-of-the-art accuracy on multiple computer vision tasks, including image classification and object detection. State-of-the-art models such as DINOv2-G [25] have over 1 billion parameters. It is important to address the efficiency of these models when deploying on edge devices.
There are many existing model acceleration techniques for ViT’s such as quantization [18], knowledge distillation [39], low-rank factorization [6], and optimized kernels [7] for attention operations. These methods often aim to reduce the complexity of the transformer attention mechanism, or learn a compressed representation that has similar effectiveness to a higher-performing model or representation.
One approach to reduce the latency of transformer models is token sparsification, removing less-important tokens [1, 15, 31, 30] to reduce workload size. These methods are easily applied to a variety of ViT architectures [34, 25, 10]. A potential benefit of sparsification approaches is that they do not necessarily require any training or fine-tuning [36, 1, 2] . Some methods [30, 42, 5, 37] may require significant training time, e.g. 100+ epochs, to recover more accuracy and yield latency reductions. Training poses a high barrier to application; training-free methods are more accessible [1, 36].
2.2 The GPU Tail Effect
The relationship between latency and workload size can be non-linear as seen in Fig. 1. Large changes in latency are caused by the GPU Tail Effect [40], arising due to workload partitioning and dispatching to the cores of a device’s GPU [23]. If, at any layer, the GPU cannot process the entire workload at once another wave of computation will be launched. A partially full wave costs the same as a full wave. The overhead introduced by a partial wave accumulates on a layer-wise basis. Consequently, a minor change in workload size may lead to disproportionately large change in latency since the overhead of partial waves is reduced across all layers. For example, on the NVIDIA AGX Orin, removing one token can change latency by up to 33% (Fig. 1).
Prior work has exploited the tail effect to guide efficient methods for machine learning models [41, 40], particularly in the context of convolutional neural network (CNN) channel pruning [13]. These methods often reduce latency with negligible accuracy degradation. However, ViT token pruning mechanisms are fundamentally different from CNN channel pruning approaches due to architectural differences between CNN’s and ViT’s. Therefore, we hypothesize that the GPU tail effect can lower latency for ViT’s, and recognize that new techniques will be needed. Our method validates this hypothesis.
3 Token Pruning with Consideration of the GPU Tail Effect
In Sec. 2.1 and Sec. 2.2, we discussed the convenience factor of training-free token pruning methods and how previous work leveraged the GPU tail effect for efficiency benefits. Therefore, we choose two design goals for our method: (1) require no training or fine-tuning of a pre-trained ViT model; and (2) achieve more and/or better accuracy-latency tradeoffs by pruning tokens according to the GPU tail effect. Tab. 1 illustrates qualitative differences between our work and others as a result of these goals.
Method Training Free? Hardware Aware?
Ours
ToMe [1]
Zero-TPrune [36]
Top-K [12],[16]
DynamicViT [30]
STViT [5]
TPS [37]
Table 1: Qualitative differences between our pruning approach and similar techniques.
As depicted in Fig. 2, our token pruning approach consists of two main parts. First, we determine a pruning schedule for a particular model and device pair — this includes identifying the number of tokens to remove and the layers at which to remove them. Second, we develop a training-free pruning technique to remove non-informative tokens during inference time. In Sec. 3.1 we show how to decide the number of tokens to prune based on the GPU tail effect. Next, in Sec. 3.2 we explain our choice of which layers to prune, completing our offline pruning schedule selection. Finally, Sec. 3.3 describes our token pruning mechanism which is used at inference time.
Figure 2: Illustration of our method to decide a pruning schedule (left) and how we prune according to the schedule at inference time (right).
3.1 Deciding a Number of Tokens to Prune for our Pruning Schedule
Workload size definition. Before explaining how to leverage the GPU tail effect, we first define workload size. We model ViT workload size as being proportional to the number of input tokens $n$, embedding dimension $d$, and batch size $b$ in Equation (1). This is because the runtime complexity of the self-attention mechanism is $O(n^{2}\times d)$ [35]; thus the workload size is proportional to $n$ and $d$. The batch size $b$ is simply a multiplier for workload size. Transformer models are (partially) parameterized by a particular choice of $d$. For DeiT-S, $d$ is $384$; for DinoV2-G, $d$ is $1536$. Fig. 1 illustrates the workload size vs latency relationships by varying $n$, $d$, and $b$ across a set of devices. Eq. 1 describes the relationship in terms of the same quantities:
$\text{Workload Size}\propto n\times d\times b.$ (1)
We measure and leverage the GPU tail effect’s characteristics for a particular workload size and device to expand the set of (better) accuracy-latency tradeoffs for token pruning.
Measuring the GPU tail effect rather than predicting it. Prior work has modeled GPU tail effect behavior for a convolutional neural network with some assumptions [17]: (1) The latency staircase pattern increases monotonically; (2) The latency difference between steps is uniform. Our experiments indicate that these assumptions may not hold for certain workload sizes using transformer networks.
First, model latency may actually increase as workload size decreases for certain workload sizes. We demonstrate this later in Sec. 4. Fig. 1 (a) demonstrates this phenomenon for DeiT-B on the TX2. This anomaly was due to PyTorch [27] launching a different set of kernels that had higher latency for one set of workload sizes.
Second, the latency difference between steps can be nonuniform. The distance between steps varies across workload sizes as seen in Fig. 1 (a-c). Considering these observations, we empirically measure the GPU tail effect due to the difficulty in predicting or modeling its behavior. The next part describes the necessary measurements.
Assigning utility to tokens for deciding a pruning schedule. Given a pre-trained model $M$ with a number of input tokens $N$, we frame selection of the number of tokens to prune $R$ as a multi-objective optimization problem between latency gains and accuracy degradation. This is the offline computation related to selecting our pruning schedule for $M$ on a target device, as seen in Fig. 2.
First, we measure latency $L(n)$ and accuracy degradation $A(n)$ of $M$ on the target device for each number of tokens to keep $n\in[1,2,...,N]$. Measuring latency of $M$ is performed with a grid search across $n$. For estimating accuracy degradation, we need a heuristic that does not depend on our pruning schedule selection. Furthermore it is useful to underestimate the accuracy of $M$ so our selection algorithm is hesitant to remove too many tokens, which can degrade accuracy significantly. A simple proxy for the accuracy of each $n$ is to apply random token removal after the first layer on $M$ [26]. It is assumed that any token pruning method should be better than random token removal since random token pruning does not consider token information content at all. Furthermore, pruning at the first layer will degrade accuracy more than pruning later in the network. Therefore random token pruning is a suitable heuristic for estimating accuracy degradation.
[…]
$\text{U}_{\text{A}}(n)=\frac{A(n)}{\text{max}\,A(n)}$ (2)
$\text{U}_{\text{L}}(n)=1-\frac{L(n)}{\text{max}\,L(n)}$ (3)
[…]
Fig. 3 illustrates an example of the plots of $L(n)$, $A(n)$ and the overall utility $U(n)$. With $N=257$, we follow Eq. 4 to obtain $N_{\text{keep}}=91$ and $R=257-91=166$. As part of our evaluation, we measure the effect of different values of $\alpha$ (Sec. 4.2).
Figure 3: Example offline computation to decide the number of tokens prune. We find $n=N_{\text{keep}}$ with maximum utility and by extension the number of tokens to prune $R$ per Eq. 4. Configuration: $\alpha$=0.5, $M$=DinoV2-G, and AGX Orin.
3.2 Pruning Schedule: Deciding Layers at which to Prune
Now we explain the layers to which our pruning mechanism will be applied. This is the final step of the offline computation illustrated in Fig. 2.
Our schedule prunes all $R$ tokens at one layer, fairly early in the network. This differs from other methods such as ToMe [1], Top-K [12, 16], and DynamicViT [30] that progressively prune tokens from layer to layer. Our design is based on two observations supported by our evaluation. First, on small, edge-grade workloads, repeated pruning operations themselves can introduce non-trivial latency (Sec. 4.5). Second, the GPU Tail Effect will lead to improved latency for each layer after which pruning has occurred — earlier pruning means more layers will benefit (Sec. 4.3).
Fig. 4 illustrates the difference between single-layer and progressive pruning. Our approach attempts to place the token pruning at or near the first 25% of layers, similar to the first pruning layer of DynamicViT [30]. For DinoV2-G, we prune at the 10th layer, and for DeiT models, we prune at the 4th layer. This allows us to realize latency reductions according to the GPU tail effect for the remaining 30 layers in DinoV2-G and the 8 remaining layers for DeiT-S and DeiT-B. We discuss the effect of pruning location in Sec. 4.3.
Figure 4: Illustration of the difference between our pruning mechanism and Token Merging [1] ToMe increases latency of DeiT-S for single batch inference on a Jetson TX2 [24] due to the overhead of pruning at every layer. Our method is able to yield a latency reduction by pruning fewer tokens at once earlier in the network.
3.3 Token Pruning Method
The schedule selection described in Secs. 3.1 and 3.2 identifies the location and number of tokens to prune. Now we give a training-free token pruning mechanism to decide which tokens to prune. This pruning mechanism is applied during inference (Fig. 2). The pruning schedule, computed offline, is an input to our token pruning mechanism.
In terms of design goals, we want our pruning mechanism to require (1) no training or finetuning on $M$, and (2) to introduce minimal overhead. First, requiring no fine-tuning or training is a convenience for practitioners, and allows pre-trained models to be deployed readily. Second, keeping overhead to a minimum for small workload sizes is necessary to yield latency reductions as previously mentioned in 3.2.
The primary design choice of pruning techniques is choosing a method to rank the importance of tokens. Similarly to other pruning techniques [3, 14], we rank the importance of each token $i$ by measuring the quantity of attention received from all other tokens. This is analogous to summing along the columns of the attention matrix in a transformer encoder layer.
Our method incorporates two aspects of importance, one from the $A_{\text{H}}$ matrix and one from the $V$ matrix. For attention, we proceed as follows. Recall that the core attention mechanism of transformers [35] is parameterized by the $Q$, $K$, and $V$ matrices. The attention matrix (including softmax) of a transformer encoder layer with $N$ tokens and $H$ heads is given by $A_{\text{H}}\in\mathbb{R}^{H\times N\times N}$  [35]. To convert the softmax attention matrix, $A_{\text{H}}$, into tokenwise importance scores we apply a reduction along the along the head dimension, $H$. We find the maximum along the head dimension for $A_{\text{H}}$. We found that applying max marginally increased accuracy compared to mean or sum. This reduction yields a new matrix $A\in\mathbb{R}^{N\times N}$. Now, summing across rows of $A$ yields the quantity of attention each token $i$ receives, similar to a component of EViT’s pruning mechanism [16] or Top-K [12]. This yields an importance score $\text{A}_{\text{S}}[i]\in\mathbb{R}^{N}$ for each token $i$.
$V_{\text{S}}$ Feature Choice Feature ImageNet1K Accuracy (%) Q 79.74 K 79.77 V 79.81
$V_{\text{S}}$ Head Reduction Operator Feature ImageNet1K Accuracy (%) mean 79.72 sum 79.70 max 79.81
Table 2: Token pruning design ablations using for our $V_{S}$ term our schedule on DeiT-S-dist for batch-size 2. In our code we use torch.amax [29] instead of max to handle the batch dimension.
Figure 5: Illustration of our token pruning method for a single transformer encoder layer at inference time. We compute an attention score and V score for each token to estimate importance, then prune the least informative tokens. Pruned tokens are aggregated into a single ‘garbage bin’ or inattentive token to retain information content as in EViT [16].
We then consider a second importance term based on the $V\in\mathbb{R}^{H\times N\times C}$ matrix. This term assigns higher scores to tokens whose entry in the $V$ matrix has a large magnitude compared with other tokens. Adding this term can marginally improve accuracy. Similarly to the attention matrix, we find the maximum along the $H$ dimension [29] — we do a set of ablations summarized in Tab. 2. Post reduction, we have $V^{\prime}\in\mathbb{R}^{N\times C}$ and we have:
$\text{V}_{\text{S}}[i]$ $=\text{softmax}(\sum^{C}_{c=1}V^{\prime}[i,c])\in\mathbb{R}^{N}.$ (5)
We add both of these scores to yield our importance score for each token $i$:
$\text{Importance}[i]$ $=\text{A}_{\text{S}}[i]+\text{V}_{\text{S}}[i].$ (6)
Finally, we borrow from EViT [16] and instead of discarding the pruned tokens, we create a new “inattentive” token based on the averaged features of all pruned tokens and append it to the set of kept tokens. This allows information from pruned tokens to be preserved in this inattentive token, which EViT found improved accuracy. Fig. 5 illustrates the token pruning process for a ViT encoder block or layer. Note that we do not prune specialized tokens such as the CLS [9] or DIST [33] tokens. These tokens serve additional purposes beyond representing the input.
4 Evaluation
In this section, we demonstrate the effectiveness of our token pruning schedule and pruning mechanism across devices (listed in Tab. 3) and across various workload sizes. To justify our design decisions, we investigate our $\alpha$ hyperparameter and pruning location choice for our schedule selection and pruning mechanism in Sec. 4.2 and Sec. 4.3, respectively. Then, we compare with a variety of existing token removal methods such as DynamicViT [30], ToMe [1]. We use Top-K [16, 12] as a baseline on ImageNet1K [8]. Sec. 4.4 demonstrates how our method advances the accuracy vs. latency pareto front across devices, workload sizes, and models. Finally, we identify some pitfalls of existing token pruning techniques for small workload sizes in Sec. 4.5. Particularly we show that some existing methods actually increase latency rather than reduce it.
4.1 Experimental Setup
Evaluation is done using three devices with varying hardware characteristics (Tab. 3). These devices feature various generations of NVIDIA GPU architectures, and thus feature different levels of computing power. We use NVIDIA TX2 [24] and NVIDIA AGX Orin [21] development boards which are edge devices designed for machine learning. Additionally, evaluation is done on a server-grade system with an NVIDIA A100 GPU [22] to demonstrate our method can generalize across a wide range of hardware, even though we design our method for edge devices. Experiments performed on the TX2 and AGX Orin were performed with fixed CPU and GPU clock rates for reproducibility and consistency. The A100 system clocks were unable to be locked because the servers machines are university-wide resources.
Device RAM (GB) GPU Cores CUDA Cores Max Power (W)
TX2 [24] 8 2 256 15
AGX Orin [21] 32 14 1792 40
A100 [22]  ✝ 80 108 6192 300
Table 3: Summary of device hardware information. We evaluate our method on two edge devices (TX2 and AGX Orin) and one server-grade system in this work.  ✝ The A100 power consumption is the power consumption only for the GPU, not the entire system.
In order to measure latency for evaluation, we use the PyTorch benchmark module [28]. All accuracy measurements were performed on the A100 system using a batch size of 512. Conveniently, accuracy only needs to be measured once for a particular pre-trained model $M$. It is only necessary to record the latency $L(n)$ on a target device for a particular workload size in order to capture GPU tail effect behavior. In subsequent experiments, we decide the pruning hyperparameters of each token pruning method for fair comparison with our method.
For reproducibility, our code is open source. We include our method, and modifications to the original ToMe codebase in order to support DinoV2 pre-trained models.
4.2 Pruning schedule ablation study across $\alpha$ (utility weight of latency/accuracy)
In 3.1 we introduce an algorithm to decide a number of tokens to prune $R$ according to the GPU tail effect. The hyperparameter of this algorithm is $\alpha$, which determines the relative weighting of the accuracy and latency utilities. Setting to $\alpha=1$ prioritizes accuracy, and only a few tokens might be pruned. Setting to $\alpha=0$ prioritizes latency, and would prune all or nearly-all tokens. To understand the performance spectrum, we perform an ablation study. The data is in Tab. 4 and illustrated in Fig. 6.
Device $\alpha$ Range Computed $R$ Accuracy Median
(# tok. to remove) Degradation Latency (ms)
AGX Orin 0.1-0.3 139 1.07 112.2 (-27.8%)
0.4-0.5 166 2.04 104.2 (-33.0%)
0.6-0.9 193 4.82 100.9 (-35.1%)
A100 0.1-0.4 73 0.25 28.77 (-4.70%)
0.5-0.9 139 1.07 28.27 (-6.34%)
Table 4: A companion table for Fig. 6 that shows the accuracy vs latency tradeoffs for the computed number of tokens to prune $R$ for various $\alpha$. Latency percent change is computed with respect to the unmodified DinoV2-G inference time.
Figure 6: Computed pruning rate $R$ from our offline pruning schedule versus various $\alpha$. Note that the selection of $R$ is relatively invariant to changes in $\alpha$. Configuration: $M$ = DinoV2-G and batch size = 2.
Evaluation on the AGX Orin for $\alpha>0.5$ yielded $R=193$ (over 75% of tokens pruned), which degraded accuracy by $\sim$2.8for 2.1% less latency. We consider this to be an unfavorable tradeoff. This ablation, in addition to results in Sec. 4.4, suggest that $\alpha\leq 0.5$ is a good choice. Thus we use $\alpha=0.5$ for all evaluations that appear in this work. Intuitively speaking this means accuracy degradation and latency reduction are considered with equal weight according to Eq. 4 when selecting $R$.
4.3 Pruning location ablation study
In 3.2 we decide at which layer our pruning mechanism should be applied. To provide insight into the potential pruning locations, we performed an ablation study. Tab. 5 illustrates latency and accuracy tradeoffs for various pruning locations of DinoV2-G.
Batch Size Pruning Accuracy Median
Layer Degradation Latency (ms)
1 1/40 3.19 68.4
10/40 1.07 81.3
20/40 0.58 93.4
30/40 0.49 104.4
2 1/40 7.08 79.1
10/40 2.04 104.7
20/40 0.97 133.0
30/40 0.83 160.8
Table 5: Latency/accuracy tradeoffs by pruning location. Configuration: $M$ = DinoV2-G on AGX Orin.
As expected, pruning earlier yields lower latency but greater accuracy degradation. For the batch size 1, our method pruned $\sim$54of input tokens at the first layer degraded accuracy by 3.19% but yielded a 40% overall latency reduction. Across both batch size 1 and 2 in this ablation study, pruning after the first 25% of layers (layer 10) results in a good balance between latency reduction and accuracy degradation. Pruning later in the network will reduce accuracy degradation, however we prioritize yielding latency benefits with our method. Therefore, we perform pruning at the layer 25% of the way into the network for all models evaluated in this work. That is layer 3 for DeiT-B and DeiT-S, and layer 10 for DinoV2-G.
4.4 Accuracy-latency tradeoffs of our method
In this section, we discuss the results of our training-free token pruning mechanism when combined with the pruning schedule inspired by the GPU tail effect. We primarily compare against ToMe [1], and include Top-K as a baseline [12, 16]. DynamicViT [30] is included in select evaluations, however unlike the other compared work and our method it requires training. Tab. 6 illustrates our method’s ability to retain higher accuracy for similar latency across various devices and workload sizes. Fig. 7 demonstrates that our token pruning mechanism and schedule expands the accuracy-latency tradeoffs on the pareto front across various workload sizes and devices.
First, our method is able to achieve higher accuracy than other training-free pruning techniques ToMe and Top-K. Tab. 6 is an ablation study across various workload sizes (batch size, models) and devices, where we tune the hyperparameters of methods such that similar latency is achieved. Unsurprisingly, DynamicViT is able to retain higher accuracy than all methods, since it requires training.
Device Batch Model and Accuracy Median
Size Method Degradation Latency (ms)
TX2 2 DeiT-B + Top-K 2.44 143.0 (-33.5%)
DeiT-B + ToMe 1.63 142.7 (-33.7%)
DeiT-B + Ours 1.16 146.0 (-32.1%)
DeiT-B + DyViT ✝ 0.57 143.7 (-33.2%)
DeiT-B 215.0 -(00.0%)
TX2 2 DeiT-S + Top-K 4.30 50.32 (-27.0%)
DeiT-S + ToMe 2.47 52.98 (-23.1%)
DeiT-S + Ours 1.24 49.44 (-28.3%)
DeiT-S + DyViT ✝ 0.46 50.88 (-26.2%)
DeiT-S 68.92 -(00.0%)
AGX Orin 2 DinoV2-G + Top-K 45.66 104.4 (-32.9%)
DinoV2-G + ToMe 6.96 104.4 (-32.9%)
DinoV2-G + Ours 2.04 104.1 (-33.1%)
DinoV2-G 155.5 -(00.0%)
A100 4 DinoV2-G + Top-K 13.31 33.76 (-16.7%)
DinoV2-G + ToMe 6.96 33.79 (-16.6%)
DinoV2-G + Ours 2.04 32.37 (-20.1%)
DinoV2-G 40.53 -(00.0%)
Table 6: Comparison with existing methods for similar latency across device and workload sizes. We tune the pruning hyperparameters of the compared methods so they achieve similar latency to our method. ✝DynamicViT (DyViT) requires training. Other evaluated methods are training-free.
Compared with ToMe and Top-K however, it is shown that our method yields consistently lower accuracy degradation. Batch size 2 inference on the AGX Orin demonstrates that to achieve similar latency to our method, Top-K degrades accuracy by more than 45.6%, while we degrade accuracy by only 2%. Across all workload sizes and devices, ToMe featured 0.47%-5.92% lower accuracy than our method (Tab. 6).
Figure 7: Illustration of accuracy and latency tradeoffs of surveyed methods with $M$ = DinoV2-G. Our pruning schedule selection pruning mechanism generate points that expand the pareto front. This indicates that considering the GPU tail effect can provide strictly better tradeoffs than existing methods in some cases. Note that the number of tokens removed at each layer ($r$) of Top-K and ToMe is evaluated from $r=5$ to $=8$ in increments of 1.
Second, we note that for larger pruning rates such as $r=7$ and $r=8$, ToMe and Top-K remove nearly all input tokens by the last layer of DinoV2-G. These high pruning ratios yield high accuracy degradation of over 40% in the case of A100 batch-size 4 for ToMe, as seen in Fig. 7. Comparatively, by pruning 54%-75% of tokens at the 10th layer of DinoV2-G according to the tail effect, we simultaneously achieve higher accuracy and lower latency than progressive pruning approaches. This shows aggressive pruning at early layers can be strictly better than progressive pruning methods for certain workload sizes. Our token pruning mechanism and schedule offers more accuracy latency tradeoffs across workload sizes.
4.5 Low workload size observations
We demonstrate there are some cases where pruning techniques actually increase latency with respect to the baseline model, rather than reduce it. Tab. 7 demonstrates an example on the TX2 for single batch inference. ToMe increased latency by $\sim$30%, while we reduced latency by $\sim$9%. Our tradeoff of 0.99% accuracy degradation and 9% latency reduction is better than ToMe and Top-K, both of which degrade accuracy and provide no latency benefit compared with the baseline DeiT-S pre-trained model. For certain workload sizes, using a baseline model or a hardware aware method such as ours is strictly better than attempting to use a token removal method with high overhead.
Device Batch Model and Accuracy Median
Size Method Degradation Latency (ms)
TX2 1 DeiT-S + Top-K (r=6) 0.73 43.38 (+18.6%)
DeiT-S + ToMe (r=6) 0.27 50.70 (+30.3%)
DeiT-S + Ours (R=69) 0.99 32.12 (-9.06%)
DeiT-S 35.32 (-0.00%)
Table 7: Given comparable total pruned tokens (69 to 72), for certain workload sizes a token pruning mechanism may increase latency due to the overhead of performing pruning at multiple layers. Our pruning method is applied at a single layer to avoid this relatively significant overhead.
We found that the operations required to do token pruning add a non-trivial latency overhead for small workload sizes on some devices such as the TX2. The overhead from masking or removing tokens can outweigh the latency reduction from processing fewer tokens. As explained in Sec. 3.2, our choice to perform single-layer pruning rather than progressive pruning allows us to yield a latency reduction.
5 Conclusion
In this work, we investigated how to improve the efficiency of pre-trained vision transformers when deployed on edge devices and provided three primary contributions. First, we provided insight into how to determine a token pruning schedule for transformer models by leveraging information about the GPU tail effect. Second, a token pruning approach was introduced that is training-free and has low overhead compared with other pruning methods. We show that our method advances the accuracy-latency pareto front compared with existing token pruning approaches across workload sizes and devices. Furthermore we show that for single batch inference, other methods may actually increase latency by 18.6-30.3% with respect to baseline, while we can reduce it by 9%. For similar latency (within 5.2%) across devices we achieve 78.6%-84.5% ImageNet1K accuracy, while the state-of-the-art, Token Merging, achieves 45.8%-85.4%.
6 Limitations
Our approach to token pruning is designed to achieve better tradeoffs at small workload sizes than other methods. This is ideal when inputs are streamed to the model, or if only small batch inference is possible. For larger workload sizes, such as running DinoV2-G with batch sizes of 32 or higher, our method may not be ideal.
7 Future Work
We demonstrate that single layer pruning can be effective even for earlier pruning layers. Learning based approaches may be able to take advantage of the high latency reductions of aggressive, early pruning while maintaining accuracy. Another promising direction is investigating the benefits of composing token sparsification and other acceleration methods together.