### v2

# Pruning One More Token is Enough: Leveraging Latency-Workload Non-Linearities for Vision Transformers on the Edge

Abstract
This paper investigates how to efficiently deploy vision transformers on edge devices. Recent methods reduce the latency of transformer neural networks by removing or merging tokens, with small accuracy degradation. However, these methods are not designed with edge device deployment in mind: jthey do not leverage information about the latency vs. workload trends to improve efficiency. First, we show the latency-workload size relationship is nonlinear for certain workload sizes. We consider this relationship to create a token pruning schedule. Second, we demonstrate a training-free, token pruning method utilizing this schedule. We show that for single batch inference, other methods increase latency by 18.6-30.3% with respect to baseline, while we can reduce it by 9%. For similar latency (within 5.2%) across devices we achieve 78.6%-84.5% ImageNet1K accuracy, while the state-of-the-art, Token Merging, achieves 45.8%-85.4%.
1 Introduction
In the past decade, Internet of Things (IoT) and commodity edge devices have become ubiquitous [4, 22, 43]. Edge devices have become sufficiently powerful, and model miniaturization techniques sufficiently capable, that machine learning (ML) models can be deployed to the network edge for various tasks, including computer vision applications [12, 38]. However, state of the art performance on various computer vision tasks are often claimed by large vision transformer (ViT) based neural network [10] architectures. ViT models such as DeiT [39] and DINOv2 [30] are not designed nor miniaturized for edge deployment. Additionally, latency is often of critical importance on edge devices, and only small inputs or workloads can be processed [18].
Prior work has shown that ViT’s have high redundancy that can be exploited for latency reduction benefits [23]. One general approach is to identify and remove parts of the input (tokens) with low information content. This is called token sparsification. Training-free token sparsification methods such as Token Merging (ToMe) [1] have been effective at reducing latency of pre-trained models. Other approaches like DynamicViT [36] can yield better accuracy than training-free methods, but require training on server-grade hardware.
[…]
We address three shortcomings of existing techniques. (1) Many existing efficient methods do not consider fine-grained hardware or latency characteristics [1, 19, 42].  Fig. 1 demonstrates the diversity of latency-workload trends across devices and workload sizes. As a result, there is room to optimize efficient methods in a hardware-aware manner by considering this relationship. (2) Some existing efficient methods may require extensive training [47, 36, 5], hindering the deployment of pre-trained models on edge devices. (3) Prior work has investigated hardware-aware methods for CNN’s [46, 45], but there is little work that shows how to handle or leverage latency-workload behavior for ViT’s. Finally, these works lack primary measurements of underlying GPU or kernel behavior. Thus, our work focuses on reducing ViT latency by considering ViT latency-workload relationships, and without requiring training. This paper has the following contributions:
  1. 1.
    We identify and profile factors that can affect ViT latency-workload relationships.
  2. 2.
    We introduce the first method using latency-workload relationships for deciding ViT token pruning schedules.
  3. 3.
    We develop a novel training-free token pruning mechanism. For similar latency (within 5.2%) across various hardware and workload sizes, we achieve 0.46 to 43.7 percentage points higher accuracy than ToMe, a state-of-the-art method.
2 Background and Related Work
In this section, we review related work on model acceleration and efficient methods for vision transformers (ViT). Some post-processing or fine-tuning methods such as quantization [21] are compatible with token sparsification techniques such as our method.
2.1 Model Acceleration of Vision Transformers
ViT [10] architectures such as DINOv2 [30] have achieved state-of-the-art accuracy on multiple computer vision tasks, including image classification and object detection. State-of-the-art models such as DINOv2-G [30] have over 1 billion parameters. It is important to address the efficiency of ViT models when deploying on edge devices. There are many existing model acceleration techniques for ViT’s such as quantization [21], knowledge distillation [44], low-rank factorization [6], and optimized kernels for attention [8]. In general, these techniques either remove redundant information or reduce network layer compute.
One approach related to redundant information removal is token pruning (sparsification) [1, 17, 37, 36]. These methods are easily applied to a variety of ViT architectures [39, 30, 10]. A potential benefit of pruning approaches is that they do not necessarily require any training or fine-tuning [41, 1, 2]. However, some methods [36, 47, 5, 42] may require significant training time, e.g. 100+ epochs, to recover more accuracy and yield latency reductions. Training poses a high barrier to application; training-free methods are more accessible [1, 41].
2.2 Latency-Workload Relationship
The thesis of token pruning is that by decreasing workload (tokens), we can reduce latency. However, the relationship between latency and workload size can be non-linear as seen in Fig. 1. This relationship can stem from how workloads are dispatched to the GPU by an ML framework, framework overhead [11], and kernel design decisions [27].  Tab. 1 illustrates primary causes of kernel inefficiency.
# Cause Diagnostic Metrics
1 Launch Config Occupancy, Grid Size
2 Memory Usage Memory Bandwidth
3 Instruction Usage Math Pipe Stalls, Instructions
Table 1: Primary causes for kernel inefficiency, with associated metrics [27, 16]. Note these causes can co-occur.
An example of Cause 1 occurs where a kernel grid size is chosen such that a partial wave of computation must be launched on the GPU — this can lead to a phenomenon termed the GPU tail effect [28]. A partially full wave will incur the same latency cost as a full wave, and this can accumulate on a layer-wise basis. Consequently, a minor change in workload size may lead to a disproportionately large change in latency, since the overhead of partial waves accumulates across layers. For example, on the NVIDIA AGX Orin [24], removing one token can decrease latency by up to 33% (Fig. 1.b).
There are many factors at play that affect latency and kernel behavior. Hardware differences, version-to-version ML framework changes, and the use of closed-source backend libraries like cudNN [7] and cuBLAS [26] make it difficult to model or predict neural network latency behavior. To illustrate this difficulty, we show more examples in Sec. 3.1.
Prior work has exploited the tail effect, which is related to Cause 1, to guide pruning methods for convolutional networks [46, 45]. Kernel-based optimization methods such as FlashAttention [8] may attempt to choose kernel launch configurations that maximize occupancy. Quantization techniques address Cause 2 by using data types that require fewer bits than 32-bit floating point for network parameters. Cause 3 can be addressed by choosing low-level operators that might be faster at the cost of precision, or vice-versa [27].
In this work, we address token pruning in the context of ViT models. Previous work frames CNN channel pruning in the context of Cause-1 problems, specifically the GPU tail effect. However, ViT token pruning mechanisms are fundamentally different from previous CNN channel pruning [14] approaches due to architectural differences between CNN’s and ViT’s. We hypothesize and later demonstrate that latency-workload relationships can also be leveraged to make better token pruning decisions for ViT’s.
3 Token Pruning with Consideration of Latency and Workload Size
In Sec. 2.1 and Sec. 2.2, we discussed the convenience of training-free token pruning methods and how previous work considered latency and workload size relationships for efficiency benefits. Therefore, we set two design goals: (1) require no training or fine-tuning of a pre-trained ViT model; and (2) achieve more and/or better accuracy-latency tradeoffs by pruning tokens according to these relationships. Tab. 2 illustrates qualitative differences between our work and others as a result of these goals.
Method Training Free Hardware Aware
Ours
ToMe [1]
EViT [19]
Top-K [13],[19]
DynamicViT [36]
TPS [42]
Table 2: Qualitative differences between our pruning approach and similar techniques. We consider ViT latency-workload relationships in our pruning, making us hardware aware.
As depicted in Fig. 3, our token pruning approach consists of two main parts. First, we determine a pruning schedule for a particular model and device pair — this includes identifying the number of tokens to remove and the layers at which to remove them. Second, we develop a training-free pruning technique to remove non-informative tokens during inference time. In Sec. 3.1 we show how to decide the number of tokens to prune based on the latency-. Next, in Sec. 3.2 we explain our choice of which layers to prune, completing our offline pruning schedule selection. Finally, Sec. 3.3 describes our token pruning mechanism which is used at inference time.
Figure 2: Latency-workload characteristics of attention operators in PyTorch [34]. Flash, MemEfficient, and Math are optimized, while Vanilla is not. Median latency was measured over 100 runs for each token count. All measurements had an IQR $<$ 1us. The two annotated latency changes of MemEff are discussed in Sec. 3.1.
Figure 3: Illustration of our method to decide a pruning schedule (left), and how we prune according to the schedule at inference time (right).
3.1 Deciding a Number of Tokens to Prune for our Pruning Schedule
ViT workload size. Before explaining our method we discuss the core operator of ViT models, the attention mechanism [40]. ViT models are parameterized (partially) by an embedding dimension $d$, and their inputs have batch size $b$ and a number of tokens $n$. The input size, or workload, for attention is a tensor with shape $b\times n\times d$. Fig. 1 illustrates the workload size vs latency relationships by varying $n$, $d$, and $b$ across a set of devices.
This wide variety of behavior across devices and workload sizes leads us to consider how ViT latency-workload relationships can be reasonably measured or modeled. We now explain underlying reasons for latency-workload non-linearity, and how we decide to measure it.
Measuring latency behavior effects vs. predicting them. Inm As previously mentioned, the GPU tail effect is one phenomenon that arises due to suboptimal kernel grid size choice. Prior work has modeled [20] or utilized [46, 45] GPU tail effect behavior with respect to convolutional neural networks CNN’s. Other work has noted the (sometimes drastic) effect of latency overhead from ML frameworks [11].
Yu et al. [45] notes that the latency of CNN’s increases in discrete jumps, and in all other workload size intervals latency remains the same. In our experiments, we find that internal ML framework operator selection can lead to a range of latency behavior. This claim for CNN’s does not hold in our evaluations for ViT’s. Fig. 2 visually illustrates the latency-workload characteristics of various attention operators in PyTorch [32] with $b$=1 and varying $n$ for a DinoV2-G [30] sized attention operation. This was measured on an NVIDIA RTX 3090 Ti GPU. Additionally, we find that the tail effect is not always the primary factor for drastic latency changes, as was described in previous work. We illustrate these findings with measurements of metrics mentioned in Sec. 2.2 and their co-occurrence with latency behavior.
In one example from Fig. 2, the MemEff attention operator features a $\sim$40latency increase from 128 to 130 tokens that is correlated with a $\sim$40increase in pipeline stall or wait time (Cause 3). But, the kernel grid size remained the same, indicating that the tail effect is not the primary cause of this latency increase (Cause 1). A second example is the $\sim$13latency decrease of MemEff from 132 to 134 tokens is correlated with a 125% increase in kernel grid size (Cause 1) and a 34% decrease in pipeline stall time (Cause 2). These effects seem to be downstream of a different kernel being chosen due to internal heuristics. Importantly, we note latency can even decrease as workload increases, due to underlying kernel, hardware, and framework behaviors.
Considering these observations, we empirically measure the latency-workload relationship due to the difficulty in predicting or modeling its behavior. The next section describes how we perform this measurement.
Ranking token importance. Given a model $M$ with a number of input tokens $N$, we frame choosing the number of tokens to prune $R$ as a multi-objective optimization problem between latency gains and accuracy degradation. This is the offline computation for selecting a pruning schedule for $M$ on a target device, as seen in Fig. 3.
First, we measure latency $L(n)$ and accuracy degradation $A(n)$ of $M$ on the target device for each number of tokens to keep $n\in[1,2,...,N]$. Measuring latency and accuracy of $M$ is performed with a grid search across $n$. We note that in some cases this can be expensive — however we find in many cases it is sufficient to perform the grid search with a stride $>1$ to save time. $A(n)$ only needs to be measured once for a particular $M$, and does not need to be done on the target device. It is only necessary to record latency $L(n)$ on the target device.
For accuracy degradation, we need to define $A(n)$ in a way that does not depend on our pruning schedule selection. Furthermore it is useful to underestimate the accuracy of $M$ so our selection algorithm is hesitant to remove too many tokens, which can degrade accuracy significantly. A simple proxy for the accuracy of each $n$ is to apply random token removal after the first layer on $M$ [31]. It is assumed that any token pruning method should be better than random token removal since random token pruning does not consider token information content at all [42]. Furthermore, pruning at the first layer will degrade accuracy more than pruning later in the network [3]. Therefore random token pruning is a suitable heuristic for estimating accuracy degradation.
Figure 4: Example offline computation to decide the number of tokens prune. We find $n=N_{\text{keep}}$ with maximum utility and by extension the number of tokens to prune $R$ per Eq. 3. Configuration: $\alpha$=0.5, $M$=DinoV2-G, and AGX Orin.
[…]
$\text{U}_{\text{A}}(n)$ $=\cfrac{A(n)}{\text{max}\,[\,A(n)\,]}$ (1)
$\text{U}_{\text{L}}(n)$ $=1-\cfrac{L(n)}{\text{max}\,[\,L(n)\,]}$ (2)
[…]
3.2 Pruning Schedule: Deciding Layers at which to Prune
Now we explain the layers to which our pruning mechanism will be applied. This is the final step of the offline computation illustrated in Fig. 3. Our schedule prunes all $R$ tokens at one layer, early in the network. This differs from other methods such as ToMe [1], Top-K [13, 19], and DynamicViT [36] that progressively prune tokens. Our design is based on two observations supported by our evaluation: First, on small workloads, repeated pruning operations themselves can introduce non-trivial latency (Sec. 4.5). Second, latency benefits compound with each layer after pruning occurs — earlier pruning means more layers will benefit (Sec. 4.3).
Compared with progressive pruning methods such as ToMe [1] that prune at every layer, our approach attempts to place the token pruning at or near the first 25% of layers, similar to the first pruning layer of DynamicViT [36]. For DinoV2-G, we prune at the 10th layer, and for DeiT models, we prune at the 4th layer. This allows us to realize latency reductions according to the GPU tail effect for the remaining 30 layers in DinoV2-G and the 8 remaining layers for DeiT-S and DeiT-B. We discuss the effect of pruning location in Sec. 4.3. As an example of the overall process,  Fig. 4 illustrates an example of the plots of $L(n)$, $A(n)$ and the overall utility $U(n)$. With $N=257$, we follow Eq. 3 to obtain $N_{\text{keep}}=91$ and $R=257-91=166$.
3.3 Token Pruning Method
The schedule selection described in Secs. 3.1 and 3.2 identifies the location and number of tokens to prune. Now we give a training-free token pruning mechanism to decide which tokens to prune. This pruning mechanism is applied during inference (Fig. 3). The pruning schedule, computed offline, is an input to our token pruning mechanism.
In terms of design goals, we want our pruning mechanism to require (1) no finetuning on $M$ and (2) to introduce minimal overhead. First, requiring no fine-tuning or training is a convenience for practitioners, and allows pre-trained models to be deployed readily. Second, keeping overhead to a minimum for small workload sizes is necessary to yield latency reductions as previously mentioned in 3.2.
The primary design choice of pruning techniques is choosing a method to rank the importance of tokens. Similarly to other pruning techniques [3, 15], we rank the importance of each token by measuring the quantity of attention received from all other tokens.
Figure 5: Illustration of our token pruning method for a single transformer encoder layer at inference time. We compute an attention score and V score for each token to estimate importance, then prune the least informative tokens. Pruned tokens are aggregated into a single ‘garbage bin’ or inattentive token to retain information content as in EViT [19].
Our method incorporates two aspects of importance, one from the $A_{\text{H}}$ matrix and one from the $V$ matrix. For the $A_{\text{H}}$ term, we proceed as follows. Recall that the core attention mechanism of transformers [40] is parameterized by the $Q$, $K$, and $V$ matrices. The attention matrix (including softmax) of a transformer encoder layer with $N$ tokens and $H$ heads is given by $A_{\text{H}}\in\mathbb{R}^{H\times N\times N}$  [40]. To convert the softmax attention matrix $A_{\text{H}}$ into tokenwise scores we: (1) reduce the head dimension $H$ with amax [35] (2) sum along the row $N$ dimension (3) normalize scores by dividing by the max. This transforms $A_{\text{H}}$ into tokenwise scores $\text{A}_{\text{S}}\in\mathbb{R}^{N}$, similar to a component of EViT’s pruning mechanism [19] or Top-K [13].
We then consider a second importance term based on the $V\in\mathbb{R}^{H\times N\times C}$ matrix. This term assigns higher scores to tokens who have high weighting in the $V$ matrix, which marginally increased accuracy. We perform a similar operation for this term: (1) reduce the $H$ dimension of $V$ with amax [35] (2) sum along the feature dimension $C$ (3) normalize scores with softmax. The resulting score vector is $V_{\text{S}}\in\mathbb{R}^{N}$, which we add to $\text{A}_{\text{S}}$ for our overall score:
Token Importance $=\text{A}_{\text{S}}+\text{V}_{\text{S}}.$ (4)
Finally, we borrow from EViT [19] and instead of discarding the pruned tokens, we create a new “inattentive” token based on the averaged features of all pruned tokens and append it to the set of kept tokens. This allows information from pruned tokens to be preserved in this inattentive token, which EViT found improved accuracy. Fig. 5 illustrates the token pruning process for a ViT encoder block or layer.
4 Evaluation
In this section, we demonstrate the effectiveness of our token pruning schedule and pruning mechanism across devices (listed in Tab. 3) and across various workload sizes. We primarily focus on smaller workloads, since these are representative of edge deployment use-cases [18]. To justify our design decisions, we investigate our $\alpha$ hyperparameter and pruning location choice for our schedule selection and pruning mechanism in Sec. 4.2 and Sec. 4.3, respectively. Then, we compare with a variety of existing token removal methods such as DynamicViT [36], ToMe [1]. We use Top-K [19, 13] as a baseline on ImageNet1K [9]. Sec. 4.4 demonstrates how our method advances the accuracy vs. latency pareto front across devices, workload sizes, and models. Finally, we identify some pitfalls of existing token pruning techniques for small workload sizes in Sec. 4.5. Particularly we show that some existing methods actually increase latency rather than reduce it.
Device GPU CUDA Max Power
Cores Cores (W)
TX2 [29] 2 256 15
Orin [24] 14 1792 40
A100 [25]  ✝ 108 6192 300
Table 3: Summary of device hardware information. We evaluate our method on two edge devices (TX2 and AGX Orin) and one server-grade system in this work.  ✝ The A100 power consumption is the power consumption only for the GPU, not the entire system.
4.1 Experimental Setup
Evaluation is done using three devices with varying hardware characteristics (Tab. 3). We use NVIDIA TX2 [29] and NVIDIA AGX Orin [24] development boards which are edge devices designed for machine learning. Additionally, evaluation is done on a server-grade system with an NVIDIA A100 GPU [25] to demonstrate our method can generalize across a wide range of hardware, even though we design our method for edge devices. Experiments performed on the TX2 and Orin were performed with fixed CPU and GPU clock rates for reproducibility and consistency. The A100 system clocks were unable to be locked because the servers machines are university-wide resources.
In order to measure latency for evaluation, we use the PyTorch benchmark module [33]. All accuracy measurements were performed on the A100 system with batch size 512 on the evaluation split of ImageNet1K. In subsequent experiments, we decide the pruning hyperparameters of each token pruning method for fair comparison with our method. For reproducibility, our code is open source.
4.2 Pruning schedule ablation study across $\alpha$ (utility weight of latency/accuracy)
In 3.1 we introduce an algorithm to decide a number of tokens to prune $R$ according to the GPU tail effect.
Device $\alpha$ R Acc. Median
Range Loss Latency (ms)
Orin [0.1, 0.3]0.10.3[0.1,\;0.3][ 0.1 , 0.3 ] 139 1.07 (-27.8%) 112.2
[0.4, 0.5]0.40.5[0.4,\;0.5][ 0.4 , 0.5 ] 166 2.04 (-33.0%) 104.2
[0.6, 0.9]0.60.9[0.6,\;0.9][ 0.6 , 0.9 ] 193 4.82 (-35.1%) 100.9
A100 [0.1, 0.4]0.10.4[0.1,\;0.4][ 0.1 , 0.4 ] 73 0.25 (-4.70%) 28.77
[0.5, 0.9]0.50.9[0.5,\;0.9][ 0.5 , 0.9 ] 139 1.07 (-6.34%) 28.27
Table 4: Accuracy vs latency tradeoffs for the computed number of tokens to prune $R$ for various $\alpha$ of our method. Latency percent change is with respect to baseline DinoV2-G inference time.
The hyperparameter of this algorithm is $\alpha$, which determines the relative weighting of the accuracy and latency utilities. Setting to $\alpha=1$ prioritizes accuracy, and only a few tokens might be pruned. Setting to $\alpha=0$ prioritizes latency, and would prune all or nearly-all tokens. To understand the performance spectrum, we performed an ablation study. The data is in Tab. 4.
Evaluation on the AGX Orin for $\alpha>0.5$ yielded $R=193$ (over 75% of tokens pruned), which degraded accuracy by $\sim$2.8for 2.1% less latency. We consider this to be an unfavorable tradeoff. This ablation, in addition to results in Sec. 4.4, suggest that $\alpha\leq 0.5$ is a good choice. Thus we use $\alpha=0.5$ for all evaluations that appear in this work. Intuitively speaking this means accuracy degradation and latency reduction are considered with equal weight according to Eq. 3 when selecting $R$.
4.3 Pruning location ablation study
In 3.2 we decide at which layer our pruning mechanism should be applied. To provide insight into the potential pruning locations, we performed an ablation study. Tab. 5 illustrates latency and accuracy tradeoffs for various pruning locations of DinoV2-G.
Batch Pruning Acc. Median
Size Layer Loss Latency (ms)
1     01/40 3.19 68.4
10/40 1.07 81.3
20/40 0.58 93.4
30/40 0.49 104.4
2     01/40 7.08 79.1
10/40 2.04 104.7
20/40 0.97 133.0
30/40 0.83 160.8
Table 5: Latency/accuracy tradeoffs by pruning location. Configuration: $M$ = DinoV2-G on AGX Orin.
As expected, pruning earlier yields lower latency but greater accuracy degradation. For the batch size 1, our method pruned $\sim$54of input tokens at the first layer degraded accuracy by 3.19% but yielded a 40% overall latency reduction. Across both batch size 1 and 2 in this ablation study, pruning after the first 25% of layers (layer 10) results in a good balance between latency reduction and accuracy degradation. Pruning later in the network will reduce accuracy degradation, however we prioritize yielding latency benefits with our method. Therefore, we perform pruning at the layer 25% of the way into the network for all models evaluated in this work. That is layer 3 for DeiT-B and DeiT-S, and layer 10 for DinoV2-G.
Device Batch Model & Acc. Median
Size Method Loss Latency (ms)
TX2 2 DeiT-B -(00.0%) 215.0
w/ Top-K 2.44 (-33.5%) 143.0
w/ ToMe 1.63 (-33.7%) 142.7
w/ DyViT ✝ 0.57 (-33.2%) 143.7
w/ Ours 1.16 (-32.1%) 146.0
TX2 2 DeiT-S -(00.0%) 68.92
w/ Top-K 4.30 (-27.0%) 50.32
w/ ToMe 2.47 (-23.1%) 52.98
w/ DyViT ✝ 0.46 (-26.2%) 50.88
w/ Ours 1.24 (-28.3%) 49.44
Orin 2 DinoV2-G -(00.0%) 155.5
w/ Top-K 45.66 (-32.9%) 104.4
w/ ToMe 6.96 (-32.9%) 104.4
w/ Ours 2.04 (-33.1%) 104.1
A100 4 DinoV2-G -(00.0%) 40.53
w/ Top-K 13.31 (-16.7%) 33.76
w/ ToMe 6.96 (-16.6%) 33.79
w/ Ours 2.04 (-20.1%) 32.37
Table 6: Comparison with existing methods for similar latency across device and workload sizes. We tune the pruning hyperparameters of the compared methods so they achieve similar latency to our method. ✝DynamicViT (DyViT) requires training. Other evaluated methods are training-free.
4.4 Accuracy-latency tradeoffs of our method
In this section, we discuss the results of our training-free token pruning mechanism when combined with the pruning schedule inspired by the GPU tail effect. We primarily compare against ToMe [1], and include Top-K as a baseline [13, 19]. DynamicViT [36] is included in select evaluations, however unlike the other compared work and our method it requires training. Tab. 6 illustrates our method’s ability to retain higher accuracy for similar latency across various devices and workload sizes. Fig. 6 demonstrates that our token pruning mechanism and schedule expands the accuracy-latency tradeoffs on the pareto front across various workload sizes and devices.
First, our method is able to achieve higher accuracy than other training-free pruning techniques ToMe and Top-K. Tab. 6 is an ablation study across various workload sizes (batch size, models) and devices, where we tune the hyperparameters of methods such that similar latency is achieved. Unsurprisingly, DynamicViT is able to retain higher accuracy than all methods, since it requires training. Compared with ToMe and Top-K however, it is shown that our method yields consistently lower accuracy degradation. Batch size 2 inference on the AGX Orin demonstrates that to achieve similar latency to our method, Top-K degrades accuracy by more than 45.6%, while we degrade accuracy by only 2%. Across all workload sizes and devices, ToMe featured 0.47%-5.92% lower accuracy than our method (Tab. 6).
Figure 6: Illustration of accuracy-latency tradeoffs of surveyed methods with $M$ = DinoV2-G: (a) batch size=2 on AGX Orin (b) batch size=4 on A100 (c) batch size=4 on AGX Orin. Our pruning schedule and mechanism generate points that expand the pareto front. The number of tokens removed at each layer ($r$) of Top-K and ToMe is evaluated from $r=5$ to $=8$ in increments of 1.
Second, we note that for larger pruning rates such as $r=7$ and $r=8$, ToMe and Top-K remove nearly all input tokens by the last layer of DinoV2-G. These high pruning ratios yield high accuracy degradation of over 40% in the case of A100 batch-size 4 for ToMe, as seen in Fig. 6. Comparatively, by pruning 54%-75% of tokens at the 10th layer of DinoV2-G according to the tail effect, we achieve higher accuracy and lower latency than progressive pruning approaches. This shows aggressive pruning at early layers can be strictly better than progressive pruning methods for certain workload sizes.
4.5 Low workload size observations
We demonstrate there are some cases where pruning techniques actually increase latency with respect to the baseline model, rather than reduce it. Tab. 7 demonstrates an example on the TX2 for single batch inference. ToMe increased latency by $\sim$30%, while we reduced latency by $\sim$9%. Our tradeoff of 0.99% accuracy degradation and 9% latency reduction is better than ToMe and Top-K, both of which degrade accuracy and provide no latency benefit compared with the baseline DeiT-S pre-trained model. For certain workload sizes, using a baseline model or a hardware aware method such as ours is strictly better than attempting to use a token removal method with high overhead.
Device Method Acc. Median
Loss Latency (ms)
TX2 Baseline (-0.00%) 35.32
Top-K (r=6) 0.73 (+18.6%) 43.38
ToMe (r=6) 0.27 (+30.3%) 50.70
Ours (R=69) 0.99 (-9.06%) 32.12
Table 7: Experiment on DeiT-S with batch size 1 that illustrates pruning overhead. Given comparable total pruned tokens (69 to 72), for certain workload sizes a token pruning mechanism may increase latency due to the overhead of performing pruning at multiple layers. Our pruning method is applied at a single layer to avoid this relatively significant overhead.
We found that the operations required to do token pruning add a non-trivial latency overhead for small workload sizes on some devices such as the TX2. The overhead from masking or removing tokens can outweigh the latency reduction from processing fewer tokens. As explained in Sec. 3.2, our choice to perform single-layer pruning rather than progressive pruning allows us to yield a latency reduction.
5 Limitations
Our approach to token pruning is designed to achieve better tradeoffs when latency-workload is non-linear. This occurs primarily for small workload sizes and batch sizes, which are common contexts in edge deployment. For large workloads such as DinoV2-G with batch size 256+, our method is less effective because latency-workload relationships becomes linear. Last, though we avoid training, our method requires a grid-search. In some contexts, this may be expensive (many tokens, large batch size, large model). These contexts are outside the focus of this work, but may be relevant for some practitioners.
6 Conclusion
In this work, we investigated how to improve the efficiency of vision transformers by considering latency-workload relationships. We made three main contributions. First, we provide insight on how ViT latency is affected by underlying hardware, machine learning framework, and kernel launch factors. Second, we show how to determine a token pruning schedule by leveraging information about latency-workload relationships. Third, we introduced a token pruning mechanism that is training-free and low overhead. Our method advances the accuracy-latency Pareto front compared with existing token pruning approaches across workload sizes and devices. For single batch inference, other methods may increase latency by 18.6-30.3% with respect to baseline, while we reduce it by 9%. For similar latency (within 5.2%) across devices, the state-of-the-art, ToMe, achieves 45.8%-85.4% ImageNet1K accuracy, while we achieve 78.6%-84.5%. Improving kernel designs to reduce non-linear latency behavior is a promising future direction.